# Distributionally Robust Control Synthesis for Stochastic Systems with Safety and Reach-Avoid Specifications

### v1

@@ Appendix A Proof of Measurability of Stochastic Kernel
Let $\varepsilon>0$ and $F_{\varepsilon}=\{y\in\mathcal{X}|\inf_{x\in F}d(x,y)<\varepsilon\}$. For small enough $\varepsilon$, $F_{\varepsilon}^{c}$ is a closed non-empty set. Let $g_{\varepsilon}\to[0,1]$ be the uniformly continuous real valued function on $\mathcal{X}$ that equals $1$ on $F$ and 00 on $(F_{\varepsilon})^{c}$ and $g_{\varepsilon}(x)\in(0,1)$ for every $x\notin F\cup F_{\varepsilon}^{c}$ (the existence of $g_{\varepsilon}$ is justified by the Urysohn’s lemma [8, Lem 7.1]). For all $\varepsilon>0$,
@@
where both terms converge to 00 thanks to the uniform continuity of $g_{\varepsilon}$ and the weak convergence of $\mu_{n}$. Thus we have proved the continuity of $(z,\mu)\mapsto\int_{V}g_{\varepsilon}(z,w)\mathrm{d}\mu(w)$ for all $\varepsilon>0$. Thus we have
@@
Similarly, $q_{\gamma_{t}}:X\times U\ni(x,u)\mapsto\delta_{\gamma_{t}(x,u)}\in\mathcal{P}(\mathbb{D})\subseteq\mathcal{P}(\mathcal{P}(W))$ is measurable. This is due to the fact that $W$ is metrizable and seperable thus $\mathcal{P}(W)$, $\mathcal{P}(\mathcal{P}(W))$, as well as $\mathcal{P}(\mathbb{D})$ are metrizable and separable thanks to [8, prop 7.20], and $q_{\gamma_{t}}$ is the composition of $\delta:\mathbb{D}\to\mathcal{P}(\mathbb{D})$ and $\gamma_{t}$.
@@ Appendix B Proofs Omitted in Main Body
By inductive hypothesis, $\mathbf{v}_{t+1}:X\to[0,1]$ is upper semi-continuous. From [8, Prop. 7.32(b)] and Lemma 3, we know that $\dot{\mathbf{v}}_{t}(x,u)$ is upper semi-continuous w.r.t. $X\times U$. We further define $\ddot{\mathbf{v}}_{t}:X\to\mathbb{R}$ by $\ddot{\mathbf{v}}_{t}(x)=\sup_{u\in U}\dot{\mathbf{v}}_{t}(x,u)$. By [8, Prop. 7.32(a)] and upper semi-continuity of $\dot{\mathbf{v}}_{t}(x,u)$, $\ddot{\mathbf{v}}_{t}$ is upper semi-continuous. From definition of integral, we have $\forall(x,u)\in X\times U$,
$0\leq\inf_{x^{\prime}\in X}\mathbf{v}_{t+1}(x^{\prime})\leq\dot{\mathbf{v}}_{t}(x,u)\leq\sup_{x^{\prime}\in X}\mathbf{v}_{t+1}(x^{\prime})\leq 1$
@@
Thus (1) is true. Now it comes to (2). From Lemma 3, we know that $\mathbf{H}(x,u,\mu,\mathbf{v}_{t+1})$ is upper semi-continuous w.r.t. $(x,u,\mu)\in X\times U\times\mathbb{D}$ for any $t=0,1\dots,T-1$. Since $\mathbb{D}$ is a closed-subspace of the compact space $\mathcal{P}(W)$, it’s compact thus separable. From [8, Prop. 7.34], it holds that for any $\epsilon>0$, there exists a measurable function $\gamma^{\star}_{t}:X\times U\to\mathbb{D}$ such that $\forall(x,u)\in X\times U$,

### v2

@@ Appendix A Proof of Measurability of Stochastic Kernel
Let $\varepsilon>0$ and $F_{\varepsilon}=\{y\in\mathcal{X}|\inf_{x\in F}d(x,y)<\varepsilon\}$. For small enough $\varepsilon$, $F_{\varepsilon}^{c}$ is a closed non-empty set. Let $g_{\varepsilon}\to[0,1]$ be the uniformly continuous real valued function on $\mathcal{X}$ that equals $1$ on $F$ and 0 on $(F_{\varepsilon})^{c}$ and $g_{\varepsilon}(x)\in(0,1)$ for every $x\notin F\cup F_{\varepsilon}^{c}$ (the existence of $g_{\varepsilon}$ is justified by the Urysohn’s lemma [9, Lem 7.1]). For all $\varepsilon>0$,
@@
where both terms converge to 0 thanks to the uniform continuity of $g_{\varepsilon}$ and the weak convergence of $\mu_{n}$. Thus we have proved the continuity of $(z,\mu)\mapsto\int_{V}g_{\varepsilon}(z,w)\mathrm{d}\mu(w)$ for all $\varepsilon>0$. Thus we have
@@
Similarly, $q_{\gamma_{t}}:X\times U\ni(x,u)\mapsto\delta_{\gamma_{t}(x,u)}\in\mathcal{P}(\mathbb{D})\subseteq\mathcal{P}(\mathcal{P}(W))$ is measurable. This is due to the fact that $W$ is metrizable and seperable thus $\mathcal{P}(W)$, $\mathcal{P}(\mathcal{P}(W))$, as well as $\mathcal{P}(\mathbb{D})$ are metrizable and separable thanks to [9, prop 7.20], and $q_{\gamma_{t}}$ is the composition of $\delta:\mathbb{D}\to\mathcal{P}(\mathbb{D})$ and $\gamma_{t}$.
@@ Appendix B Proofs Omitted in Main Body
By inductive hypothesis, $\mathbf{v}_{t+1}:X\to[0,1]$ is upper semi-continuous. From [9, Prop. 7.32(b)] and Lemma 3, we know that $\dot{\mathbf{v}}_{t}(x,u)$ is upper semi-continuous w.r.t. $X\times U$. We further define $\ddot{\mathbf{v}}_{t}:X\to\mathbb{R}$ by $\ddot{\mathbf{v}}_{t}(x)=\sup_{u\in U}\dot{\mathbf{v}}_{t}(x,u)$. By [9, Prop. 7.32(a)] and upper semi-continuity of $\dot{\mathbf{v}}_{t}(x,u)$, $\ddot{\mathbf{v}}_{t}$ is upper semi-continuous. From definition of integral, we have $\forall(x,u)\in X\times U$,
$0\leq\inf_{x^{\prime}\in X}\mathbf{v}_{t+1}(x^{\prime})\leq\dot{\mathbf{v}}_{t}(x,u)\leq\sup_{x^{\prime}\in X}\mathbf{v}_{t+1}(x^{\prime})\leq 1.$
@@
Thus (1) is true. Now it comes to (2). From Lemma 3, we know that $\mathbf{H}(x,u,\mu,\mathbf{v}_{t+1})$ is upper semi-continuous w.r.t. $(x,u,\mu)\in X\times U\times\mathbb{D}$ for any $t=0,1\dots,T-1$. Since $\mathbb{D}$ is a closed-subspace of the compact space $\mathcal{P}(W)$, it’s compact thus separable. From [9, Prop. 7.34], it holds that for any $\epsilon>0$, there exists a measurable function $\gamma^{\star}_{t}:X\times U\to\mathbb{D}$ such that $\forall(x,u)\in X\times U$,
@@
Appendix C The Computed Barrier Certificates and Control policies in Cases 2 and 3
For case $2$ with $\theta_{1}=0.1$, the computed DR-CBC is $\tilde{\mathbf{v}}_{1}(x)$ and its associated control function is $\mathbf{u}_{1}(x)$. Similarly, the results with $\theta_{2}=0.01$ are $\tilde{\mathbf{v}}_{2}(x)$ and $\mathbf{u}_{2}(x)$.
$\tilde{\mathbf{v}}_{1}(x)$ $=-0.0002964x^{4}+0.0127x^{3}-0.1396x^{2}-0.02132x+0.9917$
$\mathbf{u}_{1}(x)$ $=1.837e^{-6}x^{4}+3.752e^{-6}x^{3}-0.002694x^{2}+0.01221x+1.888$
$\tilde{\mathbf{v}}_{2}(x)$ $=-1.691e^{-5}x^{4}+0.003356x^{3}-0.1835x^{2}-0.05924x1+0.9851$
$\mathbf{u}_{2}(x)$ $=6.812e^{-8}x^{4}-8.57e^{-6}x^{3}+0.0007303x^{2}-0.0308x+1.517$
For case $3$, the computed DR-CBC is $\tilde{\mathbf{v}}(x_{1},x_{2},x_{3},x_{4})$ and its associated control function is $\mathbf{u}(x_{1},x_{2},x_{3},x_{4})$.
$\tilde{\mathbf{v}}(x_{1},x_{2},x_{3},x_{4})=$
$-0.2498x_{1}^{4}-0.4521x_{1}^{3}x_{2}+6.218e^{-16}x_{1}^{3}x_{3}-2.981e^{-18}x_{1}^{3}x_{4}+0.03694x_{1}^{3}-0.3542x_{1}^{2}x_{2}^{2}+6.744e^{-16}x_{1}^{2}x_{2}x_{3}$
$+3.973e^{-17}x_{1}^{2}x_{2}x_{4}+0.08168x_{1}^{2}x_{2}-0.6311x_{1}^{2}x_{3}^{2}-5.627e^{-5}x_{1}^{2}x_{3}x_{4}-9.227e^{-16}x_{1}^{2}x_{3}-0.233x_{1}^{2}x_{4}^{2}+2.663e^{-18}x_{1}^{2}x_{4}$
$-0.9242x_{1}^{2}-0.3872x_{1}x_{2}^{3}+4.514e^{-16}x_{1}x_{2}^{2}x_{3}-1.482e^{-17}x_{1}x_{2}^{2}x_{4}+0.05211x_{1}x_{2}^{2}-0.2759x_{1}x_{2}x_{3}^{2}+0.0001882x_{1}x_{2}x_{3}x_{4}$
$-7.616e^{-16}x_{1}x_{2}x_{3}-0.4317x_{1}x_{2}x_{4}^{2}+4.62e^{-17}x_{1}x_{2}x_{4}+0.156x_{1}x_{2}+1.353e^{-15}x_{1}x_{3}^{3}-1.112e^{-16}x_{1}x_{3}^{2}x_{4}+0.08263x_{1}x_{3}^{2}$
$+4.98e^{-16}x_{1}x_{3}x_{4}^{2}-0.003198x_{1}x_{3}x_{4}+1.376e^{-16}x_{1}x_{3}-1.151e^{-17}x_{1}x_{4}^{3}+0.001856x_{1}x_{4}^{2}+1.674e^{-17}x_{1}x_{4}-0.0001112x_{1}$
$-0.2656x_{2}^{4}+5.129e^{-16}x_{2}^{3}x_{3}+6.286e^{-17}x_{2}^{3}x_{4}+0.006773x_{2}^{3}-0.6401x_{2}^{2}x_{3}^{2}+0.0004495x_{2}^{2}x_{3}x_{4}+2.007e^{-16}x_{2}^{2}x_{3}$
$-0.2745x_{2}^{2}x_{4}^{2}+3.15e^{-17}x_{2}^{2}x_{4}-0.9106x_{2}^{2}+6.695e^{-16}x_{2}x_{3}^{3}+9.216e^{-17}x_{2}x_{3}^{2}x_{4}+0.03181x_{2}x_{3}^{2}+5.334e^{-16}x_{2}x_{3}x_{4}^{2}$
$-0.002691x_{2}x_{3}x_{4}+1.647e^{-16}x_{2}x_{3}+6.922e^{-17}x_{2}x_{4}^{3}+0.003741x_{2}x_{4}^{2}-2.836e^{-18}x_{2}x_{4}-0.0001082x_{2}-0.5929x_{3}^{4}$
$+0.001256x_{3}^{3}x_{4}-2.266e^{-16}x_{3}^{3}-0.5173x_{3}^{2}x_{4}^{2}-1.794e^{-16}x_{3}^{2}x_{4}-0.8271x_{3}^{2}+3.418e^{-5}x_{3}x_{4}^{3}-2.969e^{-17}x_{3}x_{4}^{2}$
$+3.805e^{-5}x_{3}x_{4}+2.483e^{-18}x_{3}-0.01936x_{4}^{4}+9.653e^{-18}x_{4}^{3}-0.9872x_{4}^{2}-1.874e^{-19}x_{4}+0.9999$
$\mathbf{u}(x_{1},x_{2},x_{3},x_{4})=$
$0.01773x_{1}^{4}-0.0004027x_{1}^{3}x_{2}+3.374e^{-16}x_{1}^{3}x_{3}-1.196e^{-16}x_{1}^{3}x_{4}+0.0007027x_{1}^{3}-0.0216x_{1}^{2}x_{2}^{2}-6.616e^{-17}x_{1}^{2}x_{2}x_{3}$
$+3.953e^{-17}x_{1}^{2}x_{2}x_{4}+0.002888x_{1}^{2}x_{2}-0.08656x_{1}^{2}x_{3}^{2}-3.015e^{-5}x_{1}^{2}x_{3}x_{4}-2.731e^{-16}x_{1}^{2}x_{3}-0.02166x_{1}^{2}x_{4}^{2}-1.158e^{-16}x_{1}^{2}x_{4}$
$-0.03417x_{1}^{2}+0.003754x_{1}x_{2}^{3}-7.503e^{-16}x_{1}x_{2}^{2}x_{3}-1.241e^{-16}x_{1}x_{2}^{2}x_{4}+0.00351x_{1}x_{2}^{2}-0.002516x_{1}x_{2}x_{3}^{2}$
$+0.0009659x_{1}x_{2}x_{3}x_{4}-3.586e^{-18}x_{1}x_{2}x_{3}-0.0003268x_{1}x_{2}x_{4}^{2}-1.388e^{-16}x_{1}x_{2}x_{4}-0.005415x_{1}x_{2}-2.386e^{-16}x_{1}x_{3}^{3}$
$-1.456e^{-16}x_{1}x_{3}^{2}x_{4}+0.005017x_{1}x_{3}^{2}-7.791e^{-16}x_{1}x_{3}x_{4}^{2}-0.006335x_{1}x_{3}x_{4}-1.291e^{-15}x_{1}x_{3}-1.341e^{-16}x_{1}x_{4}^{3}$
$+0.002866x_{1}x_{4}^{2}+1.0e^{-16}x_{1}x_{4}+0.001509x_{1}+0.04643x_{2}^{4}+2.58e^{-17}x_{2}^{3}x_{3}-2.009e^{-16}x_{2}^{3}x_{4}+0.0003539x_{2}^{3}-0.02193x_{2}^{2}x_{3}^{2}$
$+0.000234x_{2}^{2}x_{3}x_{4}-2.226e^{-17}x_{2}^{2}x_{3}+0.05167x_{2}^{2}x_{4}^{2}-2.946e^{-17}x_{2}^{2}x_{4}-0.07807x_{2}^{2}+5.043e^{-17}x_{2}x_{3}^{3}+6.089e^{-17}x_{2}x_{3}^{2}x_{4}$
$+0.0002895x_{2}x_{3}^{2}-1.03e^{-16}x_{2}x_{3}x_{4}^{2}-0.001483x_{2}x_{3}x_{4}-3.494e^{-17}x_{2}x_{3}-3.811e^{-16}x_{2}x_{4}^{3}+0.0006056x_{2}x_{4}^{2}+2.91e^{-16}x_{2}x_{4}$
$-0.0001431x_{2}+0.01818x_{3}^{4}+0.0003328x_{3}^{3}x_{4}-6.936e^{-17}x_{3}^{3}-0.02228x_{3}^{2}x_{4}^{2}+9.971e^{-17}x_{3}^{2}x_{4}-0.03461x_{3}^{2}-0.0001048x_{3}x_{4}^{3}$
$+2.35e^{-17}x_{3}x_{4}^{2}-1.252e^{-5}x_{3}x_{4}+9.886e^{-17}x_{3}+0.04612x_{4}^{4}+1.458e^{-18}x_{4}^{3}-0.07739x_{4}^{2}-1.024e^{-17}x_{4}+0.9907$